A sample of Si has electron and hole mobilities of 0.13 and 0.05 m /V-s 2 respectively at
300 K. It is doped with P and Al with doping densities of 21 3 1.510 / m and
21 3 2.510 / m respectively. The conductivity of the doped Si sample at 300 K is
1
Expert's answer
2020-09-28T08:10:39-0400

As per Question,

for electron' mobility "\mu_e=0.13"

for holes

mobility "\mu_h=0.05"


Concentration of Phosphorus atom i.e.

Donor concentration "N_D=n=1.5\times10^{10}/m^3"


Concentration of Aluminum atom i.e.

Acceptor impurities concentration "N_A=p=2.5\times 10^{10}/m^3"


Conductivity

"\sigma=q(n\mu_e+p\mu_p)"

="1.6\times 10^{-19}(1.5\times10^{10}\times0.13+2.5\times10^{10}\times0.05)"


="1.6\times10^{-9}(0.195+0.125)"


="1.6\times10^{-9}\times0.320"


=0.512"\times10^{-9}"

"=5.12\times10^{-8}" "\Omega^{-1}"